OpenStudy (anonymous):

x^(5/6)-6x^(5/12)+9=0

OpenStudy (campbell_st):

basically, this is a quadratic equation \[(x^{\frac{5}{6}})^2 -6x^{\frac{5}{6}} + 9 = 0\] and I can tell you its a perfect square. hope it helps

OpenStudy (aum):

\[ \large x^{5/6}-6x^{5/12}+9=0 \\ \large (x^{5/12})^2 - 6x^{5/12}+9=0 \\ \large ~~~~t^2~~~ - ~~~~6t~~~~+~~~9 = 0 ~~~ \text{where } t = x^{5/12}\\ \large \]

OpenStudy (aum):

Solve the quadratic in t. Then put back t = x^{5/12} and solve for x.

OpenStudy (campbell_st):

oops sorry about the mistake